# Prompts to Summaries: Zero-Shot Language-Guided Video Summarization

### v1

@@ 3.1 Scene detection
To segment the input video into coherent scenes, we apply a two stages process: initial scene detection with dynamic thresholding and scene boundary refinement.
@@ 3.1.1 Initial scene detection with dynamic thresholding
Goal: Segment the video into non-overlapping, chronologically ordered scenes, where each scene captures a coherent event or action.
Approach: For each consecutive frame pair $(F_{t},F_{t+1})$, we use PySceneDetect’s contentDetector [34] to compute the mean absolute pixel–intensity difference $D_{t}$ as:
$D_{t}=\frac{1}{H\times W}\sum_{i=1}^{H}\sum_{j=1}^{W}\left|F_{t}(i,j)-F_{t+1}(i,j)\right|$ (1)
where $H$ and $W$ are the height and width of the frame, respectively, and $F_{t}(h,w)$ denotes the value at pixel location $(h,w)$ in frame $t$. A cut is declared when $D_{t}\geq\tau$, otherwise $F_{t+1}$ is merged with the current shot. Then, scenes shorter than 2 s are merged to avoid over-fragmentation. Finally, scanning the entire video thus produces $N(\tau)$ scenes (see Figure 3).
Rather than relying on a fixed threshold $\tau$ to detect scene boundaries, we dynamically determine the optimal threshold $\tau^{*}$ by evaluating a predefined range of candidate values. For each candidate threshold $\tau_{i}\in\{\tau_{\min},\tau_{min}+\Delta\tau,\dots,\tau_{\max}\}$, we compute the number of detected scenes $N(\tau_{i})$ and select the threshold that maximizes the rate of change in the number of scenes with respect to $\tau$:
$\tau^{*}=\arg\max_{\tau_{i}}-\left(\frac{N(\tau_{i}+\Delta\tau)-N(\tau_{i})}{\Delta\tau}\right).$ (2)
When applying this method (Equation 2), we observe an initial increase at low threshold values, reaching a maximum, and then a sharp decrease with the number of detected scenes. We select the optimal threshold which yields the maximum negative slope after this peak, as shown in Figure 3, capturing the point with the most significant reduction of scene count. This slope captures the moment when noise-driven fragments disappear yet true boundaries remain (Figure 3), yielding a segmentation that balances between over- and under-segmentation while ensuring temporal coherence in the detected scenes (Figure 4(a)).
@@ 3.1.2 Scene boundary refinement
Goal: To improve coherence and semantic consistency, we refine the initial scene segmentation by merging excessively short scenes (not containing sufficient information)—often caused by noise or minor visual changes—with their most similar neighbors, reducing unnecessary fragmentation.
Approach: A scene is defined as short if it contains fewer than $M=150$ frames (approximately 5 seconds, we observe that defining short segments as those under 5 seconds is not sufficient to fully prevent over-fragmentation).
The previous stage marked scene boundaries by detecting sharp intensity changes. To avoid spurious cuts, we now refine them: any short segment whose visual embedding closely matches its neighbors is merged back, ensuring temporally coherent and visually consistent boundaries. Given a short scene $S_{i}$, we compute its mean embedding:
$\mathbf{e}_{i}=\frac{1}{|S_{i}|}\sum_{\mathbf{f}\in S_{i}}\mathbf{f},$ (3)
where $\mathbf{f}$ denotes the frame-level embedding (e.g., based on DINO [35] or CLIP [36]). We then calculate cosine similarity between $\mathbf{e}_{i}$ and the mean embeddings of its adjacent scenes, $\mathbf{e}_{i-1}$ and $\mathbf{e}_{i+1}$:
$\text{sim}_{\text{prev}}=\frac{\mathbf{e}_{i}\cdot\mathbf{e}_{i-1}}{\|\mathbf{e}_{i}\|\,\|\mathbf{e}_{i-1}\|},\quad\text{sim}_{\text{next}}=\frac{\mathbf{e}_{i}\cdot\mathbf{e}_{i+1}}{\|\mathbf{e}_{i}\|\,\|\mathbf{e}_{i+1}\|}.$ (4)
@@ 3.2 Scene description generation
Goal: Generating textual description summarizing the scene’s content and the overall video.
Approach: After establishing scene boundaries (fig. 4b), we generate textual descriptions for each detected scene and the full video using a pre-trained video-language model. The prompt template used for the VideoLM is defined as:
$\mathcal{P}_{VideoLM{}}$: “Describe this video in detail”
@@
To address memory constraints of VideoLMs when processing long frame sequences, we adopt a batch-level processing strategy: the sampled frames are divided into manageable batches, a description is generated for each batch, and finally, the batch-level descriptions are aggregated/concatenated to form a complete description for the scene or the entire video. To ensure that the description aggregation was performed correctly across the batches, we refined the generated description for each batch by enforcing consistent format. Specifically, each batch description was changed to begin with "The video continues" and ends with "The scene concludes", rather than "The video begins/ends", unless the batch occurred at the start or the end of the video (or the scene). This ensured coherence and continues description after aggregating or stitching different batch descriptions.
Capturing larger context. To enable providing the VideoLM a larger window (context) within each batch text description, we sample frames at a rate of 1 frame per second, selecting the middle frame of each second to avoid capturing transition frames or black frames that may appear at the edges of cuts. This strategy ensures the selection of representative content from each second of the video and enables the VideoLM to parse batches representing larger windows and thus larger context from the video, but are of the same size/budget. This strategy proved effective in our experiments, consistently capturing broader context, minimizing misinformation, and achieving near-zero omission of important details across videos of varying durations.
@@ 3.3 Scene level scoring
Goal: For every scene, we wish to generate a score which represent the importance of this scene as part of the whole video, based on the text descriptions as shown in Figure 4(c).
@@
Approach: We utilize the capabilities of LLMs to estimate scene-level importance scores based on the textual descriptions of both the individual scenes and the overall video, which were generated earlier. Specifically, we employ LLMs as judges to assess the contribution of each scene via a single carefully designed prompt template (shown in Section A of the appendix); For each scene, we construct an LLM query by plugging the scene description, and the overall video description, into the prompt template, as detailed in Algorithm 2 in the Appendix.
@@
The carefully designed prompt: The prompt in Listing (Appednix A) was meticulously engineered to elicit consistent, fine-grained importance scores from LLMs. First, we cast the LLM in the explicit role of an “objective evaluator”, which anchors the model’s behavior and reduces response variance. Next, we supply both the global (video-level) and local (scene-level) descriptions, ensuring the model reasons in context rather than in isolation. A five-band rubric—complete with numeric intervals, textual labels, and concrete definitions—calibrates the 1-to-100 scale, so identical semantic cues map to comparable scores across videos. Crucially, the prompt enforces a sparse-high prior: it instructs the model to reserve scores above 80 for scenes that directly advance the core narrative, while encouraging low scores for ancillary content. This bias guards against score inflation and aligns the LLM’s output with human summarization objectives, where only a handful of pivotal moments should survive into a concise synopsis. Finally, explicit reminders to “prioritize” essential scenes and “omit” secondary ones further steer the model toward selective, summary-oriented judgments—yielding reproducible importance estimates that integrate seamlessly with our downstream ranking pipeline.
3.4 User guided scoring
Goal: Guiding the summary process based on textual user-provided query.
@@ 3.4 User guided scoring
Approach. To enable user-guided video summarization, we adapt the same prompt used in the standart scene scoring stage (see Appendix A) with one key modification: we add instructions that explicitly consider the user’s query when evaluating scene importance. When a query is given, the prompt guides the language model to assess how well each scene aligns with the user’s content preference. Scenes that match key events or strongly relate to the query are given higher importance scores. In contrast, scenes that are irrelevant or conflict with the user’s intent are down-weighted.
@@
This lightweight modification enables zero-shot, query-aware summarization via prompt engineering alone, without changing the model or core scoring logic. A more detailed explanation, along with user-guided summarize examples, can be found in Appendix (section A).

### v2

@@ 3.1 Scene Detection
To segment the input video into coherent scenes, we apply a two stage process: initial scene detection with dynamic thresholding and scene boundary refinement.
@@ 3.1.1 Initial scene detection with dynamic thresholding
Goal:
Segment the video into non-overlapping, chronologically ordered scenes, where each scene captures a coherent event action.
Approach:
For each consecutive frame pair $(F_{t},F_{t+1})$, we use PySceneDetect’s contentDetector [37] to compute the mean absolute pixel–intensity difference $D_{t}$ as: $D_{t}=\frac{1}{H\times W}\sum_{i=1}^{H}\sum_{j=1}^{W}\left|F_{t}(i,j)-F_{t+1}(i,j)\right|$, where $H$ and $W$ are the height and width of the frame, respectively, and $F_{t}(h,w)$ denotes the value at pixel location $(h,w)$ in frame $t$. A cut is declared when $D_{t}\geq\tau$, otherwise $F_{t+1}$ is merged with the current shot. Then, scenes shorter than 2 s are merged to avoid over-fragmentation. Finally, scanning the entire video thus produces $N(\tau)$ scenes (see Figure  11(a) in the appendix).
Rather than relying on a fixed threshold to detect the boundaries of the scene, we dynamically determine the optimal threshold $\tau^{*}$ by evaluating a predefined range of candidate values. For each candidate threshold $\tau_{i}\in\{\tau_{\min},\tau_{min+\Delta\tau},\dots,\tau_{\max}\}$, we compute the number of detected scenes $N(\tau_{i})$ and select the threshold that maximizes the rate of change in the number of scenes with respect to $\tau$: $\tau^{*}=\arg\max_{\tau_{i}}-\frac{N(\tau_{i}+\Delta\tau)-N(\tau_{i})}{\Delta\tau}$, When applying this method, we observe an initial increase at low threshold values, reaching a maximum, and then a sharp decrease with the number of detected scenes. We select the optimal threshold which yields the maximum negative slope after this peak, as shown in Figure 11(a) in the appendix, capturing the point with the most significant reduction in scene count. This slope captures the moment when noise-driven fragments disappear yet true boundaries remain (see Figure 11 in the appendix), yielding a segmentation that balances between over- and under-segmentation while ensuring temporal coherence in the detected scenes (Figure 2(a)).
@@ 3.1.2 Scene boundary refinement
Goal:
To improve coherence and semantic consistency, we refine the initial scene segmentation by merging excessively short scenes (not containing sufficient information), often caused by noise or minor visual changes, with their most similar neighbors, reducing unnecessary fragmentation.
Approach:
A scene is defined as short if it contains fewer than $M=150$ frames (approximately 5 seconds, we observe that defining short segments as those under 5 seconds is not sufficient to fully prevent over-fragmentation).
The previous stage marked scene boundaries by detecting sharp intensity changes. To avoid spurious cuts, we now refine them: any short segment whose visual embedding closely matches its neighbors is merged back, ensuring temporally coherent and visually consistent boundaries. Given a short scene $S_{i}$, we compute its mean embedding: $\mathbf{e}_{i}=\frac{1}{|S_{i}|}\sum_{\mathbf{f}\in S_{i}}\mathbf{f}$, where $\mathbf{f}$ denotes the frame-level embedding (e.g., based on DINO [38] or CLIP [39]). We then calculate cosine similarity between $\mathbf{e}_{i}$ and the mean embeddings of its adjacent scenes, $\mathbf{e}_{i-1}$ and $\mathbf{e}_{i+1}$: $\text{sim}_{\text{prev}}=\frac{\mathbf{e}_{i}\cdot\mathbf{e}_{i-1}}{\|\mathbf{e}_{i}\|\,\|\mathbf{e}_{i-1}\|},\;\text{sim}_{\text{next}}=\frac{\mathbf{e}_{i}\cdot\mathbf{e}_{i+1}}{\|\mathbf{e}_{i}\|\,\|\mathbf{e}_{i+1}\|}$
@@ 3.2 Scene Description Generation
Goal:
Generating textual description summarizing the scene’s content and the overall video.
Approach:
After establishing scene boundaries (Figure 2b), we generate textual descriptions for each detected scene and the full video using a pre-trained video-language model. The prompt template used for the VideoLM is defined as:
$P_{\;VideoLM{}}$: "Describe this video in detail"
@@
Handling long videos and capturing larger context.
To address the memory constraints of VideoLMs when processing long frame sequences, we adopt a batch-level processing strategy: sampled frames are divided into manageable batches, a description is generated for each batch, and the resulting texts are concatenated to form the full description of a scene or an entire video. To ensure smooth transitions between batches, we enforce consistent phrasing, e.g., modifying intermediate batch outputs to begin with “The video continues” and end with “The scene concludes,” except for the first and last batches—thereby preserving temporal and semantic coherence. Furthermore, to provide the VideoLM with sufficiently broad temporal context within each batch, we uniformly sample frames at a rate of 1 frame per second, selecting the middle frame of each second to avoid transition or black frames that may occur near shot boundaries. This approach allows each batch to represent an extended yet coherent temporal window, effectively capturing high-level semantics while maintaining computational feasibility.
@@ 3.3 Scene level scoring
For every scene, we wish to generate a score which represent the importance of this scene as part of the whole video, based on the text descriptions as shown in Figure 2(c).
Approach:
@@
We utilize the capabilities of LLMs to estimate scene-level importance scores based on the textual descriptions of both the individual scenes and the overall video, which were generated earlier. Specifically, we employ LLMs as judges to assess the contribution of each scene via a single carefully designed prompt template (shown in Section  A.1 in the) appendix; For each scene, we construct an LLM query by plugging the scene description, and the overall video description, into the prompt template, as detailed in Algorithm  1 in the appendix.
The carefully designed prompt:
@@
The prompt in section A.1 in the appendix was meticulously engineered to elicit consistent, fine‑grained importance scores from LLMs. First, we cast the LLM in the explicit role of an “objective evaluator”, which anchors the model’s behavior and reduces response variance. Next, we supply both the global (video‑level) and local (scene‑level) descriptions, ensuring the model reasons in context rather than in isolation. A five‑band rubric—complete with numeric intervals, textual labels, and concrete definitions—calibrates the 1‑to‑100 scale, so identical semantic cues map to comparable scores across videos. Crucially, the prompt enforces a sparse‑high prior: it instructs the model to reserve scores above 80 for scenes that directly advance the core narrative, while encouraging low scores for ancillary content. This bias guards against score inflation and aligns the LLM’s output with human summarization objectives, where only a handful of pivotal moments should survive into a concise synopsis. Finally, explicit reminders to “prioritize” essential scenes and “omit” secondary ones further steer the model toward selective, summary‑oriented judgments—yielding reproducible importance estimates that integrate seamlessly with our downstream ranking pipeline.
User guided scoring
Goal:
Guiding the summary process based on textual user query.
Approach.
@@
To enable user-guided video summarization, we adapt the same prompt used in the standard scene scoring stage (see Section  A.1 in the appendix) with one key modification: we add instructions that explicitly consider the user’s query when evaluating scene importance. When a query is given, the prompt guides the language model to assess how well each scene aligns with the user’s content preference. Scenes that match key events or strongly relate to the query are given higher importance scores. In contrast, scenes that are irrelevant or conflict with the user’s intent are down-weighted.
@@
This lightweight modification enables zero-shot, query-aware summarization via prompt engineering alone, without changing the model or core scoring logic. A more detailed explanation, along with user-guided summarize examples, can be found in Appendix (Section  A.1 in the appendix)).
User-guided extension to the scene scoring prompt.